The perimeter of a rectangle is 32 feet. The length is three times the width. Finf the diagonal. Round to the nearest tenth.
Answer by nerdybill(7384)

The perimeter of a rectangle is 32 feet. The length is three times the width. Finf the diagonal. Round to the nearest tenth.
.
Let w = width
then
3w = length
.
Because we know the perimeter:
32 = 2(w + 3w)
32 = 2(4w)
16 = 4w
4 feet = w (width)
.
Length:
3w = 3(4) = 12 feet
.
Diagonal, we use the pythagorean theorem:
Let d = diagonal
then
d^2 = 4^2 + 12^2
d^2 = 16 + 144
d^2 = 160
d = 12.6 feet